4. A spring has a natural length of 18 inches and a force of 201 lbs is required to stretch and hold the spring to a length of 24 inches. What is the work required to stretch the spring from a length of 21 inches to a length of 26 inches?

1
Expert's answer
2022-05-27T14:30:03-0400
"F=k\Delta x"

"k=\dfrac{201\ lb}{24\ in-18\ in}=33.5\ lb/in"

"W=\dfrac{33.5\ lb/in(26\ in-21\ in)^2}{2}=418.75 \ lb\cdot in"